A manufacturer has

determined that m employees will produce a total of q units of

product per day, where

q = m(60 − m)

If the demand function is given by

p = −0.02q + 12

fifind the marginal-revenue product when m = 10.


1
Expert's answer
2021-11-26T12:18:03-0500

marginal revenue equation is the first derivative of the revenue function.

revenue function is given by "F \times Q"


"F=-0.002q +12"

"R=F \times q"

"R=( -0.002q + 12) \times q"

"R=-0.002q^2+12q"

But marginal revenue is given by the first derivative of the revenue function


"MR=\frac {\Delta R}{\Delta q} 0.002q^2+12q"


"MR=-0.004q+12"

but "q" is given by "q = m(60 \u2212 m)"

when "m=10"

"q=10(60-10)"


"q=600-100"


"q=500"


"MR=(-0.004 \times 500)+12"


"MR=-2 +12"


"MR=10"


MARGINAL REVENUE = "10units"